# Evaluating CrowdSplat: Perceived Level of Detail for Gaussian Crowds

### v1

@@ 2 Related Work
Figure 1: The Gaussian avatar at different LOD (#Gaussians): (a) Gold Standard–202k, (b) LOD3–50k, (c) LOD2–15k, (d) LOD1–3k.
@@ 3 Experiment Design
Stimuli: In our experiment, we aimed to evaluate the perceived LOD of 3DGS avatars by systematically varying the number of reconstructed 3D Gaussians and viewing distances. Following the approach of GaussianAvatar [7], we utilized monocular video recordings of a subject performing random movements for two minutes as training data. To reconstruct the avatar templates, we generated positional maps at different resolutions in $64\times 64$, $128\times 128$, $256\times 256$ and $512\times 512$, resulting in four levels of detail (LOD) represented by different number of 3D Gaussians: 3k, 15k, 50k and 202k, the latter being the gold standard (GS), as shown in Figure 1. An example stimulus, showing the GS and a lower LOD, can be seen in Figure 2.
@@
$\text{Percentage of Pixels}=100\times\frac{\text{New Pixel Height}}{\text{Baseline Pixel Height}}$
@@
Our stimuli comprised 60 video stimuli, each showing two 3DGS avatars side by side at equal distances/%pixels, for a duration of five seconds (see Figure 2). One avatar was rendered with 202k Gaussians (considered as the pseudo ground truth or gold standard), while the other was rendered with 50k, 15k, or 3k Gaussians. For additional validation and elimination of side bias, each comparison was repeated with the gold standard and comparison avatars swapped between the left and right positions. These comparisons were presented at the five discrete viewing distances/%pixels (100% to 40%). Each video featured one of two motion types: a cyclical jogging motion or a more complex martial arts performance. These motions were sourced from the AMASS dataset [17].
@@
We also found a significant main effect of LOD ($F(2,64)=65.60$, $\eta^{2}=0.672$, $p<0.001$). The effect size $\eta^{2}$ of 0.67 indicates that this was quite a large effect of accuracy based on the number of Gaussians used for rendering. Post-hoc analysis revealed significant differences between all three LOD levels (all $p<0.00001$), where accuracy for LOD1 with 3k Gaussians ($m=0.93$) was higher than for both LOD2 with 15k ($m=0.81$) and LOD3 with 50k ($m=0.60$), while LOD2 accuracy was also higher than for LOD3. Consistent with our expectations (Q2), participants could successfully differentiate each LOD from the Gold Standard at above chance (i.e., 0.5) rates, in particular for the lowest quality level.

### v2

@@ 2 Related Work
Figure 1: The Gaussian avatar at different LOD (#Gaussians): (a) Gold Standard–202k, (b) LOD3–50k, (c) LOD2–12k, (d) LOD1–3k.
@@ 3 Experiment Design
Stimuli: In our experiment, we aimed to evaluate the perceived LOD of 3DGS avatars by systematically varying the number of reconstructed 3D Gaussians and viewing distances. Following the approach of GaussianAvatar [7], we utilized monocular video recordings of a subject performing random movements for two minutes as training data. To reconstruct the avatar templates, we generated positional maps at different resolutions in $64\times 64$, $128\times 128$, $256\times 256$ and $512\times 512$, resulting in four levels of detail (LOD) represented by different number of 3D Gaussians: 3k, 12k, 50k and 202k, the latter being the gold standard (GS), as shown in Figure 1. An example stimulus, showing the GS and a lower LOD, can be seen in Figure 2.
@@
Percentage of Pixels=100×New Pixel HeightBaseline Pixel $\text{Percentage of Pixels}=100\times\frac{\text{New Pixel Height}}{\text{Baseline Pixel Height}}$
@@
Our stimuli comprised 60 video stimuli, each showing two 3DGS avatars side by side at equal distances/%pixels, for a duration of five seconds (see Figure 2). One avatar was rendered with 202k Gaussians (considered as the pseudo ground truth or gold standard), while the other was rendered with 50k, 12k, or 3k Gaussians. For additional validation and elimination of side bias, each comparison was repeated with the gold standard and comparison avatars swapped between the left and right positions. These comparisons were presented at the five discrete viewing distances/%pixels (100% to 40%). Each video featured one of two motion types: a cyclical jogging motion or a more complex martial arts performance. These motions were sourced from the AMASS dataset [17].
@@
We also found a significant main effect of LOD ($F(2,64)=65.60$, $\eta^{2}=0.672$, $p<0.001$). The effect size $\eta^{2}$ of 0.67 indicates that this was quite a large effect of accuracy based on the number of Gaussians used for rendering. Post-hoc analysis revealed significant differences between all three LOD levels (all $p<0.00001$), where accuracy for LOD1 with 3k Gaussians ($m=0.93$) was higher than for both LOD2 with 12k ($m=0.81$) and LOD3 with 50k ($m=0.60$), while LOD2 accuracy was also higher than for LOD3. Consistent with our expectations (Q2), participants could successfully differentiate each LOD from the Gold Standard at above chance (i.e., 0.5) rates, in particular for the lowest quality level.